# GOLLuM: Gaussian Process Optimized LLMs – Reframing LLM Finetuning through Bayesian Optimization

### v2

@@ B.1 Bayesian Optimization
$\mathbf{x}^{*}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$ (5)
@@ B.2 Sequential Decision Process
Bayesian optimization operates as a sequential decision-making process that balances exploration and exploitation. Key components include (1) a probabilistic surrogate model of the underlying objective function $f$ and (2) an acquisition function $\alpha$. The acquisition function guides the optimization process by proposing subsequent evaluation points. Common choices include expected improvement (EI 79), probability of improvement (PI 80), upper confidence bound (UCB 81, 82), and Thompson sampling (TS 83). Acquisition function selection over the points in the design space relies on the surrogate model and its predictive and uncertainty estimates. For example EI selects points that, in expectation, improve upon the current best observed value $f(\mathbf{x}_{\text{best}})$:
@@ B.3 Gaussian Processes and Marginal Likelihood Optimization
[…]
@@ B.4 Deep Kernel Gaussian Processes
The transformation $g_{\phi}$ can take the form of any neural architecture suitable for the data modality. The original paper applied the DKGP architecture on regression tasks with images using convolutional neural networks 34, while subsequent works have extended it to structured chemical domains using graph neural networks 66. As detailed in the main section, we apply this framework to textual chemical representations by using large language models (LLMs) as the deep kernel feature extractor. This allows us to incorporate both pretrained domain knowledge and task-specific adaptation within the BO loop.
@@ B.5 Large Language Models
LLMs process textual inputs by converting them into dense vector representations through a sequence of tokenization , embedding and attention-based transformations. Tokenization involves the process of splitting the input text into subword units (tokens) using a model-specific vocabulary (e.g., SentencePiece 84, Byte-Pair Encoding 85). The tokens are mapped to continuous vectors via learned embedding layers and passed through multiple self-attention layers that capture contextual relationships between tokens.
LLMs can follow different architectural designs: encoder-only (e.g., BERT 2), decoder-only (e.g., Qwen 52), and encoder-decoder (e.g., T5 49). Encoder-based models process the full input bidirectionally and are suited for classification and regression . Decoder-only models generate text autoregressively with causal masking . Encoder-decoder models combine both components and are often used for sequence-to-sequence tasks . The architecture choices impact the structure and pooling strategies used to extract unified representations from the variable-length token sequences.
@@ B.6 Parameter-efficient LLM Finetuning
One such approach is Low-Rank Adaptation (LoRA) 41, which injects trainable low-rank matrices into existing weight layers. Instead of updating a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank update of the form:
@@ C.1 Molecular Representations
Molecular representations have been extensively studied in chemistry, leading to (1) fingerprints 74, 86, 76, (2) Simplified Molecular Input Line Entry System (SMILES) strings 38, 39, (3) molecular graph-based features 87 or (4) more physics-informed descriptors derived from electronic structure calculations 88. Each of these representations encodes different aspects of molecular structure and properties making their utility task-dependent.
For example, molecular fingerprints can be effective for tasks involving structural similarity or substructure-driven properties, while quantum-derived features may be better suited for tasks involving electronic properties. For applications in BO, these representations typically require specialized kernel functions to capture relevant similarity 89.
@@ C.2 Reaction Representations
Chemical reactions, on the other hand, attach an additional layer of complexity beyond molecular representation including reaction conditions and procedural descriptions. In that sense, they present a unique challenge for machine learning due to an increased complexity and heterogeneous nature. Reaction conditions typically comprise multiple parameter types: numerical values (temperature, concentration, time), categorical variables (catalyst type, solvent choice), and detailed procedural descriptions, making their featurization challenging. Reaction representations used in ML range from simple one-hot encodings 73 to more elaborate reaction fingerprints 74, 75, 76, 44, quantum mechanical descriptors 78, 26 and learned representations 77.
We make extensive use of Differential Reaction Fingerprints (DRFPs) 44, previously shown to achieve state-of-the-art results on reaction optimization tasks compared to a variety of molecular and reaction descriptors 37. Generated by first computing circular fingerprints for each reactant and product and then taking their symmetric difference, DRFPs highlight the structural changes during the reaction while remaining computationally cheap. We input DRFPs to GP in reaction optimization tasks with available reaction SMILES. For other chemical optimization tasks we generate features through one-hot encoding of categorical variables, concatenated with numerical parameter values.
@@ C.3.2 Encoder-Based Models
ModernBERT 45: A compact and efficient embedding model trained with a retrieval objective, designed for high-speed and high-quality sentence representations.
UAE-Large 46: The Universal Alignment Embedding model, trained for multilingual and multimodal generalization with a strong emphasis on alignment across domains.
MXBAI-Embed 47: A large-scale embedding model from Mixedbread-AI, trained to preserve semantic similarity across a broad range of tasks, including code, math, and text.
@@ C.3.3 Encoder-Decoder Models
Encoder-decoder architectures, such as the T5 family 49, consist of two transformer modules: an encoder that processes the input sequence and a decoder that generates output sequences, typically in an autoregressive fashion. For embedding tasks, representations are typically extracted from the encoder side, which embeds the input text into a fixed-length latent representation. Compared to encoder-only models, encoder-decoder architectures are often pretrained with sequence-to-sequence objectives such as masked span prediction or denoising, making them well-suited for tasks involving paraphrasing, summarization, or input–output alignment.
@@
T5 (base variant) 49: A widely-used general-purpose model pretrained on a multi-task mixture of unsupervised and supervised NLP tasks. We use the encoder outputs as text embeddings.
T5Chem 43 : A domain-adapted variant of T5, finetuned on chemical tasks using the GT4SD framework. It is trained on a multitask mixture involving molecular property prediction, retrosynthesis, and chemical text modeling, making it more specialized for chemistry-related input sequences.
Instructor 48: An instruction-tuned encoder-decoder model trained on natural language–task pairs. It learns to produce embeddings guided by a task description (e.g., "Represent the reaction for similarity search"), making it suitable for alignment-sensitive downstream applications.
@@ C.3.4 Decoder-Only Models
Decoder-only architectures, exemplified by models in the GPT 3 family, generate outputs autoregressively by predicting each token conditioned on all previous ones. While traditionally used for generation tasks, these models can also produce dense representations of input text by extracting hidden states from specific tokens (e.g., the final token or special marker tokens). Decoder-only models are typically pretrained with causal language modeling objectives and operate unidirectionally, which distinguishes their contextual encoding behavior from encoder-based models.
@@
OpenAI Embeddings 54 A widely-used commercial API that provides text embeddings via proprietary transformer models. While the architectural details are not public, we assign them to decoder-style GPT family.
Qwen2-7B-Instruct 52 A large-scale instruction-tuned language model from Alibaba, based on a decoder-only architecture. We use this model in embedding mode by extracting the hidden state of the last non-padding token.
GTE-Qwen2-7B-Instruct 53 A retrieval-optimized variant of Qwen2, finetuned to produce sentence-level embeddings with improved performance on similarity and ranking tasks.
LLaMA 3–8B 50 Meta’s open LLaMA 3 model in its original instruction-tuned form, without additional adaptation for embeddings.
LLM2Vec Models 51 We also evaluate decoder-only LLMs adapted for embedding tasks using the LLM2Vec framework li2024llm2vec. These models, such as LLM2Vec–Meta-Llama-3 and LLM2Vec–Mistral-7B, are trained with masked next token prediction (MNTP) to enable bidirectional context modeling and use supervised mean pooling over selected internal layers. This adaptation allows decoder-only transformers to behave similarly to encoder models in embedding quality, while preserving their original architecture.
@@ C.4 Representations Overview
Model Arch. Pretraining Objective Pooling Chem. Dim. Source
@@
ModernBERT Encoder Retrieval contrastive CLS No 768 HF
@@
UAE-Large Encoder Alignment / Multimodal CLS No 1024 HF
@@
T5-Base Enc-Dec Masked span prediction Mean No 768 HF
@@
Instructor Enc-Dec Instruction alignment Weighted Mean Part.$\dagger$ 768 HF
@@
Qwen2-7B-Instruct Decoder Instruction tuning Last No 3584 HF
GTE-Qwen2 Decoder Contrastive retrieval Last No 3584 HF
LLM2Vec–LLaMA3 Decoder Supervised pooling Last No 4096 HF
LLM2Vec–Mistral Decoder Supervised pooling Last No 4096 HF
LLaMA 3–8B Decoder Instruction tuning Last No 4096 HF
@@ Appendix D Benchmarking Datasets
D.1 Buchwald-Hartwig reactions
We performed all initial investigations on a set of Buchwald-Hartwig (BH) reactions with the task of optimizing yield (0-100%). This dataset consists of 3955 reactions spanning across five distinct products (BH1-BH5). The data originates from a high-throughput experimentation (HTE) study published by Ahneman et al. 78. For each product, reactions were evaluated based on their percentage yield as determined by HPLC analysis. The design space is combinatorial across 15 reactants (aryl halides), 22 additives, 4 ligands, and 3 bases in DMSO solvent.
@@ D.1 Buchwald-Hartwig reactions
D.2 Additive screening
This dataset originates from a study on organic additives’ influence on the reactivity of complex Ni-catalysed reactions in a high-throughput experimentation (HTE) setup 90. It covers a wide range of screened additives (720) across four different reactions (Additives 1-4) and measures their effect on UV210 product area absorption. The challenge with traditional featurization methods in this dataset lies in the sole variability of the additive in the design space, while the other reaction parameters remain fixed. In such a setup, traditional one-hot encoding techniques would yield results similar to random search. Previous approaches report success with DRFP representation, comparing its performance to a set of molecular descriptors 37. Compared to this method, we achieve better results while representing the reaction SMILES through LLM embeddings and optimizing with our GP-guided finetuning approach.
D.3 Suzuki-Miyaura reactions
Suzuki-Miyaura dataset is a reaction optimization benchmark that includes high-throughput evaluated Suzuki-Miyaura cross-coupling reactions, originally introduced by Perera et al. 91. The dataset was generated using an automated nanomole-scale synthesis platform designed to explore large combinatorial reaction spaces efficiently. It contains 5760 reactions across varied combinations of 7 unique electrophile, 4 nucleophile, 11 ligands (plus one blank), 7 bases (plus one blank), 4 solvents and Pd(OAc)2 as precatalyst, while the optimization objective is maximizing yield.
D.4 Catalyst optimization – C2 yield
@@
D.5 Molecular optimization
We selected the benchmark datasets from Kristiadi et al. 32 to both test our model on molecular property optimization and compare directly to their approach (LAPEFT). Moreover, we use molecular SMILES as the textual representation of the data following their best practice and offering a fair comparison between the methods. The five datasets we present in this work span diverse scientific applications and optimization objectives:
  • Redox (1,407 samples): minimize redox potential for flow battery materials 92,
  • Solvation (1,407): minimize solvation energy 92,
  • Kinase (10,449): minimize docking score in kinase inhibitors 93,
  • Photoswitch (392): maximize the $\pi-\pi^{*}$ transition wavelength* in organic photoswitches 94,
  • PCE (10,000): maximize power conversion efficiency of photovoltaic materials 95.
D.6 General Chemistry Benchmarks
To move beyond reaction and molecular optimization we include three datasets from Olympus 96, spanning catalysis, crystallization, and process optimization, all of which involve continuous or mixed-variable optimization objectives. These datasets are commonly used in autonomous discovery and closed-loop optimization studies.
@@ E.2 Which LLM layers carry the most information?
Figure 7: Performance comparison of PEFT strategies across LLM architectures. We vary the proportion (10%, 25%, 50%) and location (top vs. bottom) of targeted linear layers using LoRA. Results show that top-layer finetuning consistently outperforms bottom-layer updates for encoder-only and $\ast$ models. For encoder-decoder models, performance is more consistent across layer locations. Based on these findings, we fix the default to targeting the top 25% of linear layers.
@@
Figures 7 and  8 show the results across four (we now separate the encoder-decoder models into chemistry-related–T5Chem and base T5) LLM architectures. For encoder-only, targeting the top layers consistently outperforms bottom-layer finetuning. This aligns with the well-established view that deeper layers in such models encode more abstract and domain-specific representations. Importantly, we omit bottom-layer results for decoder-only models (e.g., Qwen2-7B) due to instability and numerical issues encountered during training, likely caused by incompatible LoRA insertions in early layers, and only show results of the different ratio of LoRA adapted top layers (10% and 25%).
@@ E.3 Structure vs Fit and BO Performance
Surrogate models, which map the input $\mathbf{x}$ to the output $y$ in a probabilistic manner, come with their own inductive biases. GPs rely on a kernel function to define similarity between points. The choice of kernel encodes assumptions about smoothness, differentiability, and the geometry of the function to be modeled. For example, the Tanimoto kernel 89 might be better suited for binary fingerprint inputs, while Matérn kernels are broadly applicable to continuous Euclidean representations. In our study, we fix the surrogate kernel to Matérn 5/2 due to its balance between smoothness and flexibility, and its prevalence in chemical BO applications. All related analysis in this section is, therefore, built on the basis of this kernel. Future work may explore kernel-specific behaviors in structure/fit alignment. With the selected design space representation and the surrogate kernel, we are left with a set of assumptions about the structure of the objective function itself. If these assumptions are misspecified – e.g., if the representation induces a geometry not aligned with the kernel – then BO performance can degrade.
@@
We observe a strong correlation ($r=0.92$) between this normalized lengthscale and BO performance (Figure 2), suggesting that representations that allow the GP to maintain broader, smoother fits tend to support more successful optimization. A higher ratio indicates that the GP can generalize over broader regions while still resolving performance differences, ultimately leading to better acquisition decisions. This trend is consistent with the intuition that smoother fits – enabled by coherent, well-structured representation spaces – support more principled exploration and reduce overfitting to local noise.

### v3

@@ B.1 Bayesian optimization
$\mathbf{x}^{*}=\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x})$ (5)
@@ B.2 Sequential decision process
Bayesian optimization operates as a sequential decision-making process that balances exploration and exploitation. Key components include (1) a probabilistic surrogate model of the underlying objective function $f$ and (2) an acquisition function $\alpha$. The acquisition function guides the optimization process by proposing subsequent evaluation points. Common choices include expected improvement (EI 16), probability of improvement (PI 17), upper confidence bound (UCB 18, 97), and Thompson sampling (TS 98). Acquisition function selection over the points in the design space relies on the surrogate model and its predictive and uncertainty estimates. For example EI selects points that, in expectation, improve upon the current best observed value $f(\mathbf{x}_{\text{best}})$:
@@ B.3 Gaussian processes and marginal likelihood optimization
$k_{\text{Matérn-5/2}}(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\left(1+\frac{\sqrt{5}d}{\ell}+\frac{5d^{2}}{3\ell^{2}}\right)\exp\left(-\frac{\sqrt{5}d}{\ell}\right),$ (8)
@@ B.4 Deep kernel Gaussian processes
The transformation $g_{\phi}$ can take the form of any neural architecture suitable for the data modality. The original paper applied the DKGP architecture on regression tasks with images using convolutional neural networks 75, while subsequent works have extended it to structured chemical domains using graph neural networks 83. As detailed in the main section, we apply this framework to textual chemical representations by using large language models (LLMs) as the deep kernel feature extractor. This allows us to incorporate both pretrained domain knowledge and task-specific adaptation within the BO loop.
@@ B.5 Large language models
LLMs process textual inputs by converting them into dense vector representations through a sequence of tokenization , embedding and attention-based transformations. Tokenization involves the process of splitting the input text into subword units (tokens) using a model-specific vocabulary (e.g., SentencePiece 99, Byte-Pair Encoding 100). The tokens are mapped to continuous vectors via learned embedding layers and passed through multiple self-attention layers that capture contextual relationships between tokens.
LLMs can follow different architectural designs: encoder-only (e.g., BERT 30), decoder-only (e.g., Qwen 63), and encoder-decoder (e.g., T5 51). Encoder-based models process the full input bidirectionally and are suited for classification and regression . Decoder-only models generate text autoregressively with causal masking . Encoder-decoder models combine both components and are often used for sequence-to-sequence tasks . The architecture choices impact the structure and pooling strategies used to extract unified representations from the variable-length token sequences.
@@ B.6 Parameter-efficient LLM finetuning
One such approach is Low-Rank Adaptation (LoRA) 77, which injects trainable low-rank matrices into existing weight layers. Instead of updating a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank update of the form:
@@ C.1 Molecular Representations
Molecular representations have been extensively studied in chemistry, leading to (1) fingerprints 93, 101, 95, (2) Simplified Molecular Input Line Entry System (SMILES) strings 55, 56, (3) molecular graph-based features 102 or (4) more physics-informed descriptors derived from electronic structure calculations 103. Each of these representations encodes different aspects of molecular structure and properties making their utility task-dependent.
For example, molecular fingerprints can be effective for tasks involving structural similarity or substructure-driven properties, while quantum-derived features may be better suited for tasks involving electronic properties. For applications in BO, these representations typically require specialized kernel functions to capture relevant similarity 104.
@@ C.2 Reaction representations
Chemical reactions, on the other hand, attach an additional layer of complexity beyond molecular representation including reaction conditions and procedural descriptions. In that sense, they present a unique challenge for machine learning due to an increased complexity and heterogeneous nature. Reaction conditions typically comprise multiple parameter types: numerical values (temperature, concentration, time), categorical variables (catalyst type, solvent choice), and detailed procedural descriptions, making their featurization challenging. Reaction representations used in ML range from simple one-hot encodings 92 to more elaborate reaction fingerprints 93, 94, 95, 50, quantum mechanical descriptors 49, 20 and learned representations 96.
We make extensive use of Differential Reaction Fingerprints (DRFPs) 50, previously shown to achieve state-of-the-art results on reaction optimization tasks compared to a variety of molecular and reaction descriptors 22. Generated by first computing circular fingerprints for each reactant and product and then taking their symmetric difference, DRFPs highlight the structural changes during the reaction while remaining computationally cheap. We input DRFPs to GP in reaction optimization tasks with available reaction SMILES. For other chemical optimization tasks we generate features through one-hot encoding of categorical variables, concatenated with numerical parameter values.
@@ C.3.2 Encoder-based models
ModernBERT 57: A compact and efficient embedding model trained with a retrieval objective, designed for high-speed and high-quality sentence representations.
UAE-Large 58: The Universal Alignment Embedding model, trained for multilingual and multimodal generalization with a strong emphasis on alignment across domains.
MXBAI-Embed 59: A large-scale embedding model from Mixedbread-AI, trained to preserve semantic similarity across a broad range of tasks, including code, math, and text.
@@ C.3.3 Encoder-decoder models
Encoder-decoder architectures, such as the T5 family 51, consist of two transformer modules: an encoder that processes the input sequence and a decoder that generates output sequences, typically in an autoregressive fashion. For embedding tasks, representations are typically extracted from the encoder side, which embeds the input text into a fixed-length latent representation. Compared to encoder-only models, encoder-decoder architectures are often pretrained with sequence-to-sequence objectives such as masked span prediction or denoising, making them well-suited for tasks involving paraphrasing, summarization, or input–output alignment.
@@
T5 (base variant) 51: A widely-used general-purpose model pretrained on a multi-task mixture of unsupervised and supervised NLP tasks. We use the encoder outputs as text embeddings.
T5Chem 53 : A domain-adapted variant of T5, finetuned on chemical tasks using the GT4SD framework. It is trained on a multitask mixture involving molecular property prediction, retrosynthesis, and chemical text modeling, making it more specialized for chemistry-related input sequences.
Instructor 60: An instruction-tuned encoder-decoder model trained on natural language–task pairs. It learns to produce embeddings guided by a task description (e.g., "Represent the reaction for similarity search"), making it suitable for alignment-sensitive downstream applications.
@@ C.3.4 Decoder-only models
Decoder-only architectures, exemplified by models in the GPT 31 family, generate outputs autoregressively by predicting each token conditioned on all previous ones. While traditionally used for generation tasks, these models can also produce dense representations of input text by extracting hidden states from specific tokens (e.g., the final token or special marker tokens). Decoder-only models are typically pretrained with causal language modeling objectives and operate unidirectionally, which distinguishes their contextual encoding behavior from encoder-based models.
@@
OpenAI embeddings 52 A widely-used commercial API that provides text embeddings via proprietary transformer models. While the architectural details are not public, we assign them to decoder-style GPT family.
Qwen2-7B-Instruct 63 A large-scale instruction-tuned language model from Alibaba, based on a decoder-only architecture. We use this model in embedding mode by extracting the hidden state of the last non-padding token.
GTE-Qwen2-7B-Instruct 64 A retrieval-optimized variant of Qwen2, finetuned to produce sentence-level embeddings with improved performance on similarity and ranking tasks.
LLaMA 3–8B 61 Meta’s open LLaMA 3 model in its original instruction-tuned form, without additional adaptation for embeddings.
LLM2Vec models 62 We also evaluate decoder-only LLMs adapted for embedding tasks using the LLM2Vec framework. These models, such as LLM2Vec–Meta-Llama-3 and LLM2Vec–Mistral-7B, are trained with masked next token prediction (MNTP) to enable bidirectional context modeling and use supervised mean pooling over selected internal layers. This adaptation allows decoder-only transformers to behave similarly to encoder models in embedding quality, while preserving their original architecture.
@@ C.4 Representations overview
Model Arch. Pretrain. obj. Pool Chem. Dim. Source
@@
ModernBERT Encoder Retr. contrastive CLS No 768 HF
@@
UAE-Large Encoder Align./Multimod. CLS No 1024 HF
@@
T5-Base Enc-Dec Mask span pred. Mean No 768 HF
@@
[…]
@@
Qwen2-7B-Instruct Decoder Instr. tuning Last No 3584 HF
GTE-Qwen2 Decoder Contrastive retr. Last No 3584 HF
LLM2Vec–LLaMA3 Decoder Supervised pool. Last No 4096 HF
LLM2Vec–Mistral Decoder Supervised pool. Last No 4096 HF
LLaMA 3–8B Decoder Instr. tuning Last No 4096 HF
@@ Appendix D Benchmarking datasets
D.1 Synthetic organic chemistry
Buchwald-Hartwig reactions
We performed all initial investigations on a set of Buchwald-Hartwig (BH) reactions with the task of optimizing yield (0-100%). This dataset consists of 3955 reactions spanning across five distinct products (BH1-BH5). The data originates from a high-throughput experimentation (HTE) study published by Ahneman et al. 49. For each product, reactions were evaluated based on their percentage yield as determined by HPLC analysis. The design space is combinatorial across 15 reactants (aryl halides), 22 additives, 4 ligands, and 3 bases in DMSO solvent.
@@ D.1 Synthetic organic chemistry
Additive screening
This dataset originates from a study on organic additives’ influence on the reactivity of complex Ni-catalysed reactions in a high-throughput experimentation (HTE) setup 66. It covers a wide range of screened additives (720) across four different reactions (Additives 1-4) and measures their effect on UV210 product area absorption. The challenge with traditional featurization methods in this dataset lies in the sole variability of the additive in the design space, while the other reaction parameters remain fixed. In such a setup, traditional one-hot encoding techniques would yield results similar to random search. Previous approaches report success with DRFP representation, comparing its performance to a set of molecular descriptors 22. Compared to this method, we achieve better results while representing the reaction SMILES through LLM embeddings and optimizing with our GP-guided finetuning approach.
Suzuki-Miyaura reactions
Suzuki-Miyaura dataset is a reaction optimization benchmark that includes high-throughput evaluated Suzuki-Miyaura cross-coupling reactions, originally introduced by Perera et al. 65. The dataset was generated using an automated nanomole-scale synthesis platform designed to explore large combinatorial reaction spaces efficiently. It contains 5760 reactions across varied combinations of 7 unique electrophile, 4 nucleophile, 11 ligands (plus one blank), 7 bases (plus one blank), 4 solvents and Pd(OAc)2 as precatalyst, while the optimization objective is maximizing yield.
D.2 Molecular optimization
We selected the benchmark datasets from Kristiadi et al. 43 to both test our model on molecular property optimization and compare directly to their approach (LAPEFT). Moreover, we use molecular SMILES as the textual representation of the data following their best practice and offering a fair comparison between the methods. The five datasets we present in this work span diverse scientific applications and optimization objectives:
  • Redox (1,407 samples): minimize redox potential for flow battery materials 68,
  • Solvation (1,407): minimize solvation energy 68,
  • Kinase (10,449): minimize docking score in kinase inhibitors 69,
  • Photoswitch (392): maximize the $\pi-\pi^{*}$ transition wavelength* in organic photoswitches 70,
  • PCE (10,000): maximize power conversion efficiency of photovoltaic materials 71.
D.3 Material science and catalysis
To move beyond reaction and molecular optimization we include two datasets from Olympus 67, involving catalysis and crystallization, together with C2 yield optimization from 41, all of which involve continuous or mixed-variable optimization objectives. These datasets are commonly used in autonomous discovery and closed-loop optimization studies.
Catalyst optimization – C2 yield
@@ E.2 Which LLM layers carry the most information?
[…]
@@
Figures 8 and  9 show the results across four (we now separate the encoder-decoder models into chemistry-related–T5Chem and base T5) LLM architectures. For encoder-only, targeting the top layers consistently outperforms bottom-layer finetuning. This aligns with the well-established view that deeper layers in such models encode more abstract and domain-specific representations. Importantly, we omit bottom-layer results for decoder-only models (e.g., Qwen2-7B) due to instability and numerical issues encountered during training, likely caused by incompatible LoRA insertions in early layers, and only show results of the different ratio of LoRA adapted top layers (10% and 25%).
@@ E.3 Structure vs fit and BO performance
Surrogate models, which map the input $\mathbf{x}$ to the output $y$ in a probabilistic manner, come with their own inductive biases. GPs rely on a kernel function to define similarity between points. The choice of kernel encodes assumptions about smoothness, differentiability, and the geometry of the function to be modeled. For example, the Tanimoto kernel 104 might be better suited for binary fingerprint inputs, while Matérn kernels are broadly applicable to continuous Euclidean representations. In our study, we fix the surrogate kernel to Matérn 5/2 due to its balance between smoothness and flexibility, and its prevalence in chemical BO applications. All related analysis in this section is, therefore, built on the basis of this kernel. Future work may explore kernel-specific behaviors in structure/fit alignment. With the selected design space representation and the surrogate kernel, we are left with a set of assumptions about the structure of the objective function itself. If these assumptions are misspecified – e.g., if the representation induces a geometry not aligned with the kernel – then BO performance can degrade.
@@
We observe a strong correlation ($r=0.92$) between this normalized lengthscale and BO performance (Figure 3c), suggesting that representations that allow the GP to maintain broader, smoother fits tend to support more successful optimization. A higher ratio indicates that the GP can generalize over broader regions while still resolving performance differences, ultimately leading to better acquisition decisions. This trend is consistent with the intuition that smoother fits – enabled by coherent, well-structured representation spaces – support more principled exploration and reduce overfitting to local noise.